# m m OTr_ #1#3(#2#3)

### v2

@@ VIII Conclusions: what is really quantum is incompatibility
In essence, it appears to us that those phenomena rooted in the existence of incompatible variables, i.e., on the commutation relations involving the constant of nature $\hbar$, cannot be found in classical (indeterministic) theory. This is because in the latter we have only imposed that at any instant in time the state of a system has a finite volume $Vol(S)$, in phase space. However, in that case there is no relation whatsoever between the canonically conjugated variables appearing in phase space (e.g., position and momentum). In quantum theory, on the contrary, there is a smallest size for the volume (for each degree of freedom), i.e., $Vol(R)\geq(2\pi\hbar)^{\#(d.o.f.)}$, which establishes a relation between the (in)determinacy of position and momentum, a feature absent in classical theory. [^10] A partial classical analogue of uncertainty relations is discussed in Ref. [61], in the context of Gromov’s non-squeezing theorem in symplectic geometry. So, what really characterizes genuine quantum phenomena – at least at the kinematical level – is the Planck (reduced) constant $\hbar$. The dependence on $\hbar$ of genuine quantum features is manifest in uncertainty relations, [^11] In Ref. [1], it was shown that certain forms of uncertainty relations can also be retrieved in specific classical toy models (i.e., generalized non-contextual ontological models). However, we notice that the latter is an artificial toy model built to mimic certain features of quantum theory, such as incompatibility, while still adhering to some notion of classicality. The approach followed here, instead, is to take the actual theory of classical physics and relax its assumption of infinite precision, which leads to indeterminism without establishing any relation between the indeterminacy of conjugate variables. wave-particle duality [62], or weak values [25]. But this is the case also where it does not appear so explicitly, like in the violation of Bell inequalities which indeed require incompatible measurements [63], as well as in quantum key distribution [64], to name but a few examples.

### v3

@@ 8 Conclusions: what is truly quantum is incompatibility
This by no means wants to show that every quantum phenomenon can be explained by an indeterministic interpretation of classical physics. In essence, it appears to us that those phenomena rooted in the existence of incompatible variables, i.e., on the commutation relations involving a constant of nature that imposes a minimal indeterminacy of a state (which is circumstantially $\hbar$ in quantum theory), cannot be found in classical (indeterministic) theory. This is because in the latter we have only imposed that at any instant in time the state of a system has a finite volume $Vol(S)$, in phase space. However, in that case there is no relation whatsoever between the canonically conjugated variables appearing in phase space (e.g., position and momentum). In quantum theory, on the contrary, there is a smallest size for the volume (for each degree of freedom), i.e., $Vol(R)\geq(2\pi\hbar)^{\#(d.o.f.)}$, which establishes a relation between the (in)determinacy of position and momentum, a feature absent in classical theory. [^10] A partial classical analogue of uncertainty relations is discussed in Ref. [65], in the context of Gromov’s non-squeezing theorem in symplectic geometry. So, what really characterizes genuine quantum phenomena – at least at the kinematical level – is the Planck (reduced) constant $\hbar$. The dependence on $\hbar$ of genuine quantum features is manifest in the uncertainty relations, [^11] In Ref. [2], it was shown that certain forms of uncertainty relations can also be retrieved in specific classical toy models (i.e., generalized non-contextual ontological models). However, we notice that the latter is an artificial toy model built to mimic certain features of quantum theory, such as incompatibility, while still adhering to some notion of classicality. The approach followed here, instead, is to take the actual theory of classical physics and relax its assumption of infinite precision, which leads to indeterminism without establishing any relation between the indeterminacy of conjugate variables. wave-particle duality [66], or weak values [28]. But this is the case also where it does not appear so explicitly, like in the violation of Bell inequalities which indeed require incompatible measurements [67], as well as in quantum key distribution [68], to name but a few examples.